OpenStudy (anonymous):

Find a number that is between 7/11 and 0.75 . A. 25/44 B. 77/99 C.23/33 D.76/99 help please!

hero (hero):

Let's first change .75 to a fraction: .75 = 75/100 = 3/4 Do you agree @kira12341234 ?

hero (hero):

So now, the question becomes Find a number between 7/11 and 3/4 The easiest way to deal with this is to use equivalent fractions to re-write each fraction in a way so that the denominators of both are the same.

hero (hero):

Notice that the Least Common Denominator or LCD is the Least Common Multiple of 4 and 11. \(4 \times 11 = 44\)

hero (hero):

Multiply 7/11 by 4/4 and 3/4 by 11/11 to get 28/44 and 33/44

hero (hero):

so in essence we need to find a fraction, x, such that \(\dfrac{28}{44} < x < \dfrac{33}{44}\)

hero (hero):

From here, one approach you can take is cross multiplication

hero (hero):

We know that 25/44 won't work, so we won't bother with that.

hero (hero):

Also, let's not forget that we can reduce the fractions in the given choices.

hero (hero):

77/99 = 7/9

hero (hero):

So let's see if \(\dfrac{7}{11} < \dfrac{7}{9} < \dfrac{3}{4}\) will work

hero (hero):

Notice that if we cross multiply 7/11 < 7/9 we get 63 < 77 which is true If we cross multiply 7/9 < 3/4 we get 28 < 27 which is false so that doesn't work

hero (hero):

Next test: \(\dfrac{7}{11} < \dfrac{23}{33} < \dfrac{3}{4}\)

hero (hero):

Cross multiply 7/11 < 23/33 to get 231 < 253 which is true and then 23/33 < 3/4 to get 92 < 99 which is also true So the correct answer is 23/33